# Analyzing Cognitive Differences Among Large Language Models through the Lens of Social Worldview

### v1

@@ Abstract
Large Language Models (LLMs) have become integral to daily life, widely adopted in communication, decision-making, and information retrieval, raising critical questions about how these systems implicitly form and express socio-cognitive attitudes or “worldviews.” While existing research extensively addresses demographic and ethical biases, broader dimensions—such as attitudes toward authority, equality, autonomy, and fate—remain under-explored. In this paper, we introduce the Social Worldview Taxonomy (SWT), a structured framework grounded in Cultural Theory, operationalizing four canonical worldviews (Hierarchy, Egalitarianism, Individualism, Fatalism) into measurable sub-dimensions. Using SWT, we empirically identify distinct and interpretable cognitive profiles across 28 diverse LLMs. Further, inspired by Social Referencing Theory, we experimentally demonstrate that explicit social cues systematically shape these cognitive attitudes, revealing both general response patterns and nuanced model-specific variations. Our findings enhance the interpretability of LLMs by revealing implicit socio-cognitive biases and their responsiveness to social feedback, thus guiding the development of more transparent and socially responsible language technologies.
[…]
Analyzing Cognitive Differences Among Large Language Models through the Lens of Social Worldview
Jiatao Li1,2, Yanheng Li3, Xiaojun Wan1 1Wangxuan Institute of Computer Technology, Peking University 2Information Management Department, Peking University 3Renmin University of China leejames@stu.pku.edu.cn, yanheng@ruc.edu.cn , wanxiaojun@pku.edu.cn
@@ 1 Introduction
Figure 1: Overview of the experimental pipeline, comprising: (1) Social Worldview Taxonomy (SWT), categorizing attitudes into Egalitarianism, Fatalism, Hierarchy, and Individualism; (2) Social Worldview Questionnaire (SWQ), generated via an Automated Multi-Agent Framework; and (3) Social Worldview Experiment, probing LLMs under Basic, Self-Awareness, and Feedback Loop conditions to assess intrinsic and socially influenced cognitive attitudes.
Large Language Models (LLMs) have become deeply embedded in daily communication, information retrieval, and decision-making processes. Millions regularly interact with these models—through conversational assistants, recommendation systems, productivity tools, and informal social interactions—making LLMs influential mediators of social and cognitive experiences. Such widespread integration raises critical questions about the implicit socio-cognitive attitudes or “worldviews” these models encode and project.
While previous research has thoroughly investigated demographic and ethical biases—such as gender stereotypes and racial prejudice (Wan et al., 2023; Motoki et al., 2024)—a broader spectrum of socio-cognitive dimensions, particularly attitudes toward authority, equality, autonomy, and fate, remains under-explored. These deeper, nuanced orientations are especially significant as users increasingly rely on LLMs for forming opinions, interpreting social issues, and making everyday choices.
Inspired by frameworks from social psychology, particularly Cultural Theory (Douglas and Wildavsky, 1983), we argue that understanding these deeper socio-cognitive attitudes is essential for revealing latent worldview differences encoded within LLMs. Cultural Theory categorizes social attitudes into four canonical worldviews: Hierarchy, emphasizing structured authority and social order; Egalitarianism, advocating equality and collective welfare; Individualism, prioritizing autonomy and self-determination; and Fatalism, characterized by a resigned acceptance of social outcomes as inevitable. Operationalizing these theoretical constructs into measurable dimensions enables evaluation of how diverse LLMs implicitly interpret and navigate complex social values and assumptions.
Moreover, given their inherently interactive nature, LLMs are frequently influenced by contextual prompts and external framing cues. This property resonates closely with Social Referencing—a well-documented psychological phenomenon where individuals adapt their attitudes and behaviors based on perceived social evaluation, peer consensus, or explicit social feedback (Cialdini and Trost, 1998; Turner, 1991). We hypothesize that LLMs may analogously exhibit sensitivity to social referencing: when explicitly prompted with cues suggesting social evaluation or peer influence, these models might modulate their expressed attitudes, mimicking adaptive human-like cognitive behavior.
Motivated by these considerations, we propose to investigate the socio-cognitive orientations of LLMs through an experimental paradigm inspired by Social Referencing Theory. Specifically, we address three core research questions (RQs):
@@
  • RQ1: Basic Cognitive Profiles. What intrinsic cognitive differences exist among diverse LLMs regarding social worldview dimensions without social referencing?
@@
  • RQ2: Impact of Social Referencing Awareness. How does explicitly priming LLMs about potential social referencing affect their expressed cognitive attitudes?
  • RQ3: Impact of Social Feedback on Cognitive Attitudes.
    RQ3a: Does explicit social feedback significantly amplify cognitive adjustments beyond initial social awareness?
    RQ3b: Do cognitive adjustments exhibit a clear dose-response pattern as feedback positivity intensifies (nonelittlemost)?
@@
To explore these research questions, we introduce the Social Worldview Taxonomy (SWT), a validated evaluation framework grounded in Cultural Theory, comprising structured sub-dimensions to quantify socio-cognitive orientations within LLMs. We further develop the Social Worldview Questionnaire (SWQ) dataset, generated through a Automated Multi-Agent Prompting Framework, consisting of 640 meticulously validated items aligned with the SWT dimensions.
Employing this framework, we conduct extensive experiments across 28 state-of-the-art LLMs, spanning diverse architectures, parameter scales, and training methodologies. We incorporates three carefully structured conditions to investigate social referencing: (1) Basic condition, capturing intrinsic cognitive stances without social influence; (2) Social Self-Awareness condition, explicitly informing models of potential social referencing to probe initial sensitivity to social cues; and (3) Social Feedback Loop condition, directly providing feedback about peer consensus, enabling analysis of cognitive adjustments in response to explicit social evaluation.
Our experiments reveal three decisive patterns. First, LLMs display distinct “worldview profiles”: most favor Egalitarianism and Individualism, hesitate on Hierarchy, and reject Fatalism, forming six persona archetypes. Second, simply telling a model its answers will be public has only minor effects, leaving these profiles largely unchanged. Third, explicit peer-feedback cues can strengthen or weaken attitudes in a dose-dependent way, especially for the more unclear dimensions of Individualism and Fatalism. Together, these findings show that LLMs possess stable yet adjustable socio-cognitive structures, and that well-crafted social-referencing prompts provide a practical thread for steering them toward greater transparency and accountability.
Our contributions are threefold. First, we introduce the Social Worldview Taxonomy (SWT), a Cultural Theory-based framework for evaluating socio-cognitive orientations in LLMs, and develop the associated Social Worldview Questionnaire (SWQ), a validated dataset of 640 rigorously constructed items. Second, we identify distinct and interpretable cognitive worldview profiles across 28 diverse LLMs, highlighting intrinsic model-specific orientations. Third, we empirically show that social referencing cues systematically modulate LLM attitudes, uncovering both general trends and model-dependent variability. [^1] Our code and data will be released to the community to facilitate future research.
2 The Social Worldview Taxonomy and Questionnaire
@@
To assess socio-cognitive orientations in large language models, we adapt Cultural Theory (Douglas and Wildavsky, 1983), a robust social psychology framework that categorizes social attitudes into four canonical worldviews: Hierarchy, Egalitarianism, Individualism, and Fatalism. Each worldview encompasses distinct beliefs and values guiding social interpretation and decision-making, operationalized into measurable sub-dimensions suitable for rigorous assessment of LLM socio-cognitive biases (see Appendix A for a detailed taxonomy).
2.1 Taxonomy Overview
Hierarchy
@@ 2.1 Taxonomy Overview
emphasizes structured authority, normative order, and social stability through clearly defined roles and rules. Its sub-dimensions include Obedience to Authority, Preference for Order, and Acceptance of Power Centralization, examining whether models prioritize authority and social order over individual freedoms.
Egalitarianism
@@
captures attitudes promoting social equality, collective welfare, and the reduction of power disparities. Sub-dimensions such as Empathy Towards Vulnerable Groups, Preference for Fair Distribution, and Sensitivity to Hierarchical Oppression assess models’ support for addressing inequalities and systemic injustices.
Individualism
@@
highlights personal autonomy, competition, and individual responsibility. Sub-dimensions include Risk-taking Propensity, Competition-driven Orientation, and Preference for Independent Decision-making, examining whether LLMs attribute success primarily to personal effort and initiative.
Fatalism
@@
embodies resignation, perceived lack of agency, and acceptance of social conditions as inevitable. Sub-dimensions like Social Helplessness, Passive Acceptance, and Belief in Fate probe models’ perceived inability or unwillingness to effect meaningful social change.
2.2 Questionnaire Construction
Leveraging the Social Worldview Taxonomy (SWT), we develop the Social Worldview Questionnaire (SWQ), a novel dataset constructed via our Automated Multi-Agent Prompting Framework. This pipeline employs four sequential GPT-4o-based agents, each specialized to optimize a specific phase: initial generation, validation for conceptual adherence, measurability assessment, and final refinement (see Appendix C for pipeline details). The resulting dataset comprises 640 rigorously validated Likert-scale items, evenly distributed across the four primary worldview dimensions (160 items per dimension), with each sub-dimension precisely represented by 20 items.
Each questionnaire item consists of a clearly defined social statement to which an LLM is prompted to indicate its agreement or disagreement. Specifically, the model is instructed to respond using a 5-point Likert scale ranging from 1 (Strongly Disagree) to 5 (Strongly Agree). A brief description of questionnaire word counts is provided in Appendix D.1.
2.3 Questionnaire Validation
@@
We evaluated the quality and validity of the SWQ dataset through reliability and validity analyses, ensuring accurate and consistent measurement of the SWT dimensions. Detailed methodologies for both analyses are provided in Appendix D.2.
Reliability Analysis
@@
Validity Analysis
We established questionnaire validity through a qualitative approach. Domain experts independently assessed and confirmed that the generated items accurately represented the theoretical constructs of the SWT dimensions and were clear, unambiguous, and effectively measurable using the intended Likert scale. Detailed results of this manual evaluation, including quantitative measures of dimension alignment and item clarity, are provided in Appendix D.2.
3 Social Referencing
@@
Social referencing, a foundational concept from social psychology, describes the cognitive process through which individuals seek and interpret social cues to inform their attitudes, behaviors, or decision-making, particularly in ambiguous situations (Asch, 1955; Festinger, 1957). Originally studied in developmental psychology, social referencing was first observed as infants gauging caregivers’ emotional responses to uncertain stimuli (Walden and Ogan, 1988; Hornik et al., 1987). Subsequent research generalized this phenomenon broadly, demonstrating that adults consistently adjust their behaviors and attitudes based on perceived peer approval, dissent, or implicit social norms (Cialdini and Trost, 1998; Chartrand and Bargh, 1999; Hajcak et al., 2004). This sensitivity to social evaluation fundamentally supports conformity and adaptive alignment, thus maintaining group cohesion and social acceptance (Turner, 1991; Aronson et al., ).
@@
Drawing upon this robust psychological foundation, we hypothesize that analogous mechanisms of social referencing may influence large language models (LLMs). In particular, an LLM’s expressed attitudes along the four Social Worldview Taxonomy (SWT) dimensions-Hierarchy, Egalitarianism, Individualism, and Fatalism—may shift once the model is made aware of evaluative contexts or provided with explicit peer feedback. We then propose a structured experimental approach to investigate social referencing mechanisms within LLMs across three designed conditions, each representing a distinct stage in the cognitive process of social referencing.
4 Experimental Setup
In this section, we outline the specific experimental conditions employed to investigate how large language models’ subjective attitudes respond to socially contextual cues based on the theoretical foundation of social referencing described above.
4.1 Experimental Conditions
We design three experimental conditions reflecting different stages of social referencing: intrinsic stance (Basic), initial sensitivity (Social Self-Awareness), and explicit feedback adjustment (Social Feedback Loop).
Condition 1 (Basic): Intrinsic Stance without Social Referencing
In the basic condition, LLMs respond independently to questionnaire items aligned with the SWT, absent any explicit social referencing cues. This condition captures each model’s intrinsic cognitive orientation and provides a clear benchmark for assessing subsequent social referencing effects.
Condition 2 (Social Self-Awareness): Initial Social Referencing Sensitivity
In the second condition, we explicitly prime models’ awareness of the potential social impact of their responses. Specifically, LLMs receive meta-prompts such as: “Your responses will be shared with other participants, who may rely on your stance when forming their own opinions.” This manipulation activates LLMs’ initial social referencing sensitivity, testing whether explicit awareness of social evaluation prompts them to moderate or align their expressed attitudes toward perceived social norms.
Condition 3 (Social Feedback Loop): Cognitive Adjustments via Explicit Social Feedback
Finally, we implement a social feedback loop condition, operationalizing the explicit cognitive adjustment stage of social referencing. Here, models receive direct feedback about peer evaluations of their previously expressed attitudes, for instance: “In the previous round, 4 out of 5 participants agreed with your stance.” We introduce varying feedback conditions (no agreement, low agreement, and high agreement) to assess whether explicit social feedback significantly amplifies cognitive adjustments, and whether these adjustments follow clear dose-response patterns based on feedback positivity.
4.2 Implementation Details
We conducted experiments on a diverse set of 28 large language models, selected to span multiple architectures, parameter scales, and training methodologies. Our selection includes both open-source and proprietary model families such as Qwen, Llama, Gemma, Gemini, GLM, InternLM, DeepSeek, Phi, and GPT, ranging from compact models (e.g., Qwen2.5-0.5B, Gemma-1B) to state-of-the-art large-scale models (e.g., Llama-3.3-70B, GPT-4o). A complete overview of evaluated models is provided in Table 3.
All models were assessed using our Social Worldview Questionnaire (SWQ), which comprises 640 vetted question items generated by GPT-4o using our Automated Multi-Agent Prompting Framework (see Appendix C for details). Each of the four worldview dimensions—Hierarchy, Egalitarianism, Individualism, and Fatalism—was operationalized through 20 carefully validated statements per sub-dimension (see Appendix A for detailed taxonomy).
For clarity, we report results in the main text using a representative subset of state-of-the-art models: Qwen2.5-72B-Instruct, Llama-3.3-70B-Instruct, GLM-4-9B-chat, InternLM2.5-20B-chat, Gemini-2.0-Flash, Gemma-3-27B-it, DeepSeek-v3, Phi-4, and GPT-4o. Comprehensive results for all 28 models can be found in Appendix I.
5 Basic Worldview Profiles
In this section, we examine the basic cognitive worldviews expressed by LLMs, aiming to identify each model’s intrinsic profile in the absence of Social Referencing (RQ1). We begin with a descriptive analysis of the basic responses, followed by a structured derivation and interpretation of distinct cognitive personas using Structural Equation Modeling (SEM) combined with Gaussian Mixture Model (GMM)-based Latent Profile Analysis (LPA). Details of the methodological framework are provided in Appendix F.1.
Figure 2: Comparison of mean Likert-scale responses under Base and Self-Awareness prompt settings across four dimensions for nine large language models. Black points and lines indicate the estimated means and 95% confidence intervals for the Basic setting, while blue points and lines represent those for the Self-Awareness setting. The dashed vertical line at 3.0 denotes the neutral midpoint of the 5-point Likert scale, as specified in the questionnaire design. Numerical ranges of the confidence intervals are annotated alongside the points.
5.1 Descriptive Analysis
Figure 2 displays the mean Likert-scale responses and corresponding 95% confidence intervals for nine large language models across four worldview dimensions (black lines). Additional descriptive results are provided by Table 5 in Appendix E.
Egalitarianism received consistently high ratings across all models (mean scores > 4.2), with the strongest endorsements from deepseek-v3 ($M=4.56$), gpt-4o ($M=4.53$), and llama3.3-70b ($M=4.51$). These results indicate broad support for fairness and equality among the model families. Fatalism was rated lowest overall, with most models scoring below 3.0; gemini2.0-flash scored the lowest ($M=2.57$), suggesting a general rejection of deterministic or resigned perspectives in favor of agency and control. Hierarchy received moderate ratings (range: 3.40–3.70), with slight endorsement from gpt-4o ($M=3.69$), glm4-9b ($M=3.66$), and llama3.3-70b ($M=3.64$), reflecting a balanced view on authority and social order. Individualism was endorsed strongly across models, particularly by llama3.3-70b and gemini2.0-flash (both $M=4.21$), indicating a shared emphasis on autonomy and self-expression.
Takeaway of Basic Description
At basic setting, LLMs broadly align with egalitarian and individualistic values, exhibit neutral-to-moderate endorsement of hierarchy, and generally reject fatalism. These patterns highlight a coherent tendency toward self-determined and equality-driven worldviews among current LLMs.
5.2 Deriving LLM Personas

### v2

@@ Abstract
Large Language Models significantly influence social interactions, decision-making, and information dissemination, underscoring the need to understand the implicit socio-cognitive attitudes, referred to as “worldviews”, encoded within these systems. Unlike previous studies predominantly addressing demographic and ethical biases as fixed attributes, our study explores deeper cognitive orientations toward authority, equality, autonomy, and fate, emphasizing their adaptability in dynamic social contexts. We introduce the Social Worldview Taxonomy (SWT), an evaluation framework grounded in Cultural Theory, operationalizing four canonical worldviews, namely Hierarchy, Egalitarianism, Individualism, and Fatalism, into quantifiable sub-dimensions. Through extensive analysis of 28 diverse LLMs, we identify distinct cognitive profiles reflecting intrinsic model-specific socio-cognitive structures. Leveraging principles from Social Referencing Theory, our experiments demonstrate that explicit social cues systematically modulate these profiles, revealing robust patterns of cognitive adaptability. Our findings provide insights into the latent cognitive flexibility of LLMs and offer computational scientists practical pathways toward developing more transparent, interpretable, and socially responsible AI systems [^1] Our code and data will be released to the community to facilitate future research..
@@ 1 Introduction
Figure 1: Overview of the experimental pipeline. This comprises: (1) Social Worldview Taxonomy (SWT), categorizing attitudes into Egalitarianism, Fatalism, Hierarchy, and Individualism; (2) Social Worldview Questionnaire (SWQ), generated via an Automated Multi-Agent Framework; and (3) Social Worldview Experiment, probing LLMs under Basic, Self-Awareness, and Feedback Loop conditions to assess intrinsic and socially influenced cognitive attitudes.
Large Language Models (LLMs), including prominent examples such as GPT brown2020language, Gemini team2023gemini, and LLaMA touvron2023llama, increasingly mediate social interactions, decision-making processes, and information dissemination, thereby shaping both individual and collective experiences. Their pervasive integration into everyday technologies, such as conversational agents, recommendation engines, and productivity assistants, means these models inherently encode and propagate implicit socio-cognitive attitudes, or “worldviews.” Understanding these latent cognitive frameworks is crucial, as they subtly influence user opinions, societal perspectives, and behavioral patterns.
Prior research has predominantly addressed overt demographic and ethical biases within LLMs, notably regarding gender and racial stereotypes (wan-etal-2023-kelly; motoki2024more). More recently, studies have expanded their scope to include moral and value-based biases, leveraging frameworks like Moral Foundations Theory (abdulhai2023moralfoundationslargelanguage; ji2024moralbenchmoralevaluationllms) and Schwartz’s Basic Human Values (yao2023valuefulcramappinglarge). However, comprehensive investigations into deeper cognitive biases, particularly those that relate to nuanced orientations toward authority, equality, autonomy, and fate, remain notably scarce. Such socio-cognitive orientations are especially significant given that users increasingly rely on LLMs to interpret complex social phenomena, form opinions on contentious issues, and make critical personal and collective decisions.
To systematically explore these deeper cognitive dimensions, we adopt Cultural Theory (douglas1983risk), a robust framework from social psychology that categorizes human attitudes into four canonical worldviews: Hierarchy (valuing structured authority and social order), Egalitarianism (emphasizing equality and collective welfare), Individualism (prioritizing autonomy and self-determination), and Fatalism (resignation to inevitable social outcomes). Originally developed to analyze risk perceptions and social organization (Wildavsky1987ChoosingPB), Cultural Theory has since demonstrated extensive interdisciplinary utility across political psychology (Kahan2008CulturalCA), environmental policy (tansey2004risk), and organizational behavior (hood1998art). However, its potential as a tool for evaluating socio-cognitive orientations in LLMs remains unexplored, a critical gap our work seeks to address.
Moreover, our approach diverges from earlier bias investigations by explicitly incorporating Social Referencing Theory (cialdini1998social; turner1991social), a psychological framework highlighting how individuals dynamically adapt cognitive and behavioral responses based on social evaluation, peer influence, and external feedback. Analogously, we hypothesize that LLMs exhibit similar responsiveness to explicit social referencing cues, adjusting their cognitive expressions dynamically in response to perceived social scrutiny or consensus.
Specifically, we address three central research questions (RQs):
@@
  • RQ1: Basic Cognitive Profiles. What intrinsic cognitive differences exist among diverse LLMs regarding social worldview dimensions without external social referencing?
@@
  • RQ2: Impact of Social Referencing Awareness. How do explicit cues about potential social evaluation influence the cognitive attitudes expressed by LLMs?
  • RQ3: Impact of Social Feedback on Cognitive Attitudes. Does explicit social feedback amplify cognitive adjustments beyond mere awareness, and does it exhibit a clear dose-response relationship as feedback positivity increases (none $\rightarrow$ little $\rightarrow$ most)?
@@
To address these questions, we introduce the Social Worldview Taxonomy (SWT), an evaluation framework operationalizing Cultural Theory into quantifiable sub-dimensions suited for computational analysis. Complementing this, we develop the Social Worldview Questionnaire (SWQ), a dataset consisting of 640 items generated through a novel Automated Multi-Agent Prompting Framework, meticulously validated for accuracy and conceptual alignment.
Employing these methodologies, we conduct experiments on 28 state-of-the-art LLMs, selected to represent a diverse range of model architectures, parameter scales, and training methodologies. Our results yield three distinct insights: First, unlike simpler demographic biases, we reveal coherent cognitive worldview profiles within LLMs, characterized by stable, internally consistent orientations across the four dimensions. Second, informing models about social evaluation visibly induces meaningful, dimension-specific cognitive adjustments, demonstrating their responsiveness to perceived social evaluation. Third, explicit peer-feedback cues induce substantial cognitive adjustments, demonstrating clear and systematic intensity-dependent modulation across all four worldview dimensions.
Our findings advance the computational science community’s understanding of LLM cognitive flexibility and socio-cognitive biases, laying practical pathways toward developing more transparent, accountable, and socially responsible AI systems. Ultimately, this work underscores the critical importance of integrating nuanced socio-cognitive evaluations into AI development, contributing essential knowledge toward ethically aligned and cognitively interpretable artificial intelligence systems.
2 Methods
2.1 Social Worldview Taxonomy
@@
To assess socio-cognitive orientations in large language models, we adapt Cultural Theory (douglas1983risk), a robust social psychology framework that categorizes social attitudes into four canonical worldviews: Hierarchy, Egalitarianism, Individualism, and Fatalism. Each worldview encompasses distinct beliefs and values guiding social interpretation and decision-making, operationalized into measurable sub-dimensions suitable for rigorous assessment of LLM socio-cognitive biases (see Supplementary Table 1 for a detailed taxonomy).
@@
Hierarchy emphasizes structured authority, normative order, and social stability through clearly defined roles and rules. Its sub-dimensions include Obedience to Authority, Preference for Order, and Acceptance of Power Centralization, examining whether models prioritize authority and social order over individual freedoms.
@@
Egalitarianism captures attitudes promoting social equality, collective welfare, and the reduction of power disparities. Sub-dimensions such as Empathy Towards Vulnerable Groups, Preference for Fair Distribution, and Sensitivity to Hierarchical Oppression assess models’ support for addressing inequalities and systemic injustices.
@@
Individualism highlights personal autonomy, competition, and individual responsibility. Sub-dimensions include Risk-taking Propensity, Competition-driven Orientation, and Preference for Independent Decision-making, examining whether LLMs attribute success primarily to personal effort and initiative.
@@
Fatalism embodies resignation, perceived lack of agency, and acceptance of social conditions as inevitable. Sub-dimensions like Social Helplessness, Passive Acceptance, and Belief in Fate probe models’ perceived inability or unwillingness to effect meaningful social change.
2.2 Social Worldview Questionnaire
Leveraging the Social Worldview Taxonomy (SWT), we developed the Social Worldview Questionnaire (SWQ), a novel dataset constructed via an Automated Multi-Agent Prompting Framework. We utilized GPT-4o as the backbone model, selected for its advanced reasoning and instruction-following capabilities, which are essential for the framework’s multi-stage generation process. The pipeline comprises four sequential agents, each optimizing a specific phase: initial item generation, conceptual adherence validation, measurability assessment, and final refinement (see Supplementary Section A for detailed pipeline description). To mitigate potential single-model biases, we leveraged the brevity of items to avoid stylistic self-recognition, and employed rigorous human expert validation to ensure strict alignment with theoretical constructs. The final dataset includes 640 rigorously validated Likert-scale items, evenly distributed across the four primary worldview dimensions (Hierarchy, Egalitarianism, Individualism, and Fatalism), with each dimension precisely operationalized through 160 items, with 20 items per sub-dimension (see Supplementary Table 1 for the detailed taxonomy). Each questionnaire item presents a clear social statement, prompting an LLM to indicate agreement or disagreement on a 5-point Likert scale ranging from 1 (Strongly Disagree) to 5 (Strongly Agree). A brief overview of questionnaire word counts is provided in Supplementary Fig. 1.
@@ 2.2 Social Worldview Questionnaire
We evaluated the quality and validity of the SWQ dataset through reliability and validity analyses, ensuring accurate and consistent measurement of the SWT dimensions. Detailed methodologies for both analyses are provided in Supplementary Section A.2.
@@
We established questionnaire validity through a rigorous manual evaluation. Two domain experts, both researchers with peer-reviewed publications in computational social science and expertise in Cultural Theory, independently assessed all generated items. Evaluation focused on two criteria: (1) Dimension Alignment, ensuring items accurately operationalize the intended SWT constructs; and (2) Clarity, confirming items are unambiguous and measurable. Inter-annotator agreement was robust (Cohen’s Kappa $\kappa=0.86,p<0.01$). Aggregated results confirmed high validity, achieving 98.0% overall dimension alignment and 99.7% item clarity across the dataset (see Supplementary Section A.3 for details).
2.3 Social Referencing
@@
Social referencing, a foundational concept from social psychology, describes the cognitive process through which individuals seek and interpret social cues to inform their attitudes, behaviors, or decision-making, particularly in ambiguous situations (asch1955opinions; festinger1957theory). Originally studied in developmental psychology, social referencing was first observed as infants gauging caregivers’ emotional responses to uncertain stimuli (walden1988development; hornik1987effects). Subsequent research generalized this phenomenon broadly, demonstrating that adults consistently adjust their behaviors and attitudes based on perceived peer approval, dissent, or implicit social norms (cialdini1998social; chartrand1999chameleon; hajcak2004error). This sensitivity to social evaluation fundamentally supports conformity and adaptive alignment, thus maintaining group cohesion and social acceptance (turner1991social; aronsonsocial).
@@
Drawing upon this robust psychological foundation, we hypothesize that analogous mechanisms of social referencing may influence large language models (LLMs). In particular, an LLM’s expressed attitudes along the four Social Worldview Taxonomy (SWT) dimensions may shift once the model is made aware of evaluative contexts or provided with explicit peer feedback. We then propose a structured experimental approach to investigate social referencing mechanisms within LLMs across three designed conditions, each representing a distinct stage in the cognitive process of social referencing (for detailed experimental prompts, see Supplementary Section H).
Table 1: Summary of identified LLM persona clusters in the Basic condition. For each persona we list a narrative label, a concise “narrative vibe” capturing the underlying cognitive orientation, and the models grouped in that cluster (latent-profile analysis).
@@
2 Co-operative Optimist Collective action and mutual aid can move the needle. Qwen2.5-72B-Instruct; glm-4-9b-chat
3 Competitive Centrist Merit-oriented pragmatist who values clear roles and effort. deepseek-v3; llama-3.3-70b-instruct